### v1

# U-MARVEL: Unveiling Key Factors for Universal Multimodal Retrieval via Embedding Learning with MLLMs

Abstract
Universal multimodal retrieval (UMR), which aims to address complex retrieval tasks where both queries and candidates span diverse modalities, has been significantly advanced by the emergence of MLLMs. While state-of-the-art MLLM-based methods in the literature predominantly adopt contrastive learning principles, they often differ in their specific training recipes. Despite their success, the mechanisms underlying their retrieval capabilities remain largely unexplored, potentially resulting in suboptimal performance and limited generalization ability. To address these issues, we present a comprehensive study aimed at uncovering the key factors that drive effective embedding learning for UMR using MLLMs. We begin by implementing a general MLLM-based embedding learning pipeline, and systematically analyze the primary contributors to high-performing universal retrieval systems. Based on this, we explore various aspects of the details in embedding generation and training strategies, including progressive transition, hard negative mining and re-ranker distillation. Notably, our findings reveal that often-overlooked factors can have a substantial impact on model performance. Building on these discoveries, we introduce a unified framework termed U-MARVEL (Universal MultimodAl RetrieVal via Embedding Learning), which outperforms state-of-the-art competitors on the M-BEIR benchmark by a large margin in supervised settings, and also exihibits strong zero-shot performance on several tasks such as composed image retrieval and text-to-video retrieval. These results underscore the generalization potential of our framework across various embedding-based retrieval tasks. Code is available at https://github.com/chaxjli/U-MARVEL
1 Introduction
Multimodal information retrieval [37] has emerged as a pivotal research direction at the intersection of multimodal understanding and information retrieval. It serves as a cornerstone for a wide range of modern AI applications, including retrieval-augmented generation (RAG) [23, 10], text-image retrieval [3, 60], and visual question answering (VQA) [8, 9]. While existing methods such as CLIP [42], BLIP [28, 27], and CoCa [56] have demonstrated impressive performance in cross-modal retrieval, they often struggle to address the diverse and complex requirements encountered in real-world scenarios.
To tackle these challenges, the research community has increasingly focused on universal multimodal retrieval (UMR) [51], depicted in Fig. 1(a), which aims to develop unified, instruction-guided retrievers capable of handling diverse retrieval tasks across multiple modalities. Recent advancements in this area, such as LamRA [36], MM-Embed [30], GME [62], UniME [15] and VLM2VEC [24], have shown promising results on UMR powered by MLLMs. However, most of the existing approaches often directly adapt MLLMs to embedding tasks without systematically examining training schemes tailored for MLLM-based embedding models. Consequently, the impact of design decisions on retrieval performance remains inadequately understood, highlighting the need for a more systematic investigation—a gap that motivates our study.
In this paper, we aim to address these issues by systematically exploring the design principles for building high-performing universal multimodal retrievers with pre-trained MLLMs (Fig. 1(b)). To this end, we first implement a simple yet effective embedding learning pipeline based on contrastive learning, following the existing approaches in the literature. Subsequently, we conduct a comprehensive study to investigate three key aspects for training high-performance retrieval models built on MLLMs. Specifically, we explore: (1) how to adapt decoder-only MLLMs into instruction-guided embedding models; (2) how to train MLLM-based embedders within the context of a contrastive learning framework; and (3) how to distill the recall-then-rerank paradigm into a single model to further enhance retrieval performance. Our findings highlight several overlooked factors critical to retrieval performance, including the superiority of bidirectional attention with mean pooling over last-token mechanisms, and the interactions among parameters of batch size, learning rate, and temperature coefficient in contrastive learning. We also discover that integrating hard negative mining and knowledge distillation effectively mitigates computational inefficiencies in the recall-then-rerank pipeline, while simultaneously improving retrieval efficiency and accuracy.
Motivated by these insights, we present U-MARVEL, a unified framework for learning universal multimodal embeddings tailored to UMR tasks. Our method achieves significant improvements over state-of-the-art UMR approaches on the M-BEIR benchmark [51] in supervised settings, and also demonstrates strong zero-shot generalization abilities on tasks such as text-to-video retrieval and composed image retrieval.
The contributions of this paper can be summarized as follows. (1) We conduct a comprehensive exploration of the design space for MLLM-based universal retrieval, uncovering key factors that improve performance and providing actionable insights for future research. The findings presented in this paper aim to advance the effectiveness of UMR tasks within the research community. (2) We introduce U-MARVEL, a unified framework that achieves state-of-the-art performance in both supervised and zero-shot settings, showcasing its effectiveness across a wide range of retrieval tasks.
2 Related Work
Multimodal Large Language Models.
Large language models (LLMs) [1, 54, 32, 45], including multimodal LLMs [61, 55], have achieved remarkable success across a wide range of general-purpose tasks. For instance, LLaVA [34] integrates image and text processing for tasks like visual question answering, image description, and visual reasoning. The Qwen-VL series [47, 2] enhances this by supporting flexible image resolutions and extending capabilities to video understanding through temporal modeling and localization. InternVL3 [63] unifies multimodal and linguistic learning by leveraging both multimodal data and text corpora in a single pre-training stage.
Multimodal Retrieval.
Multimodal retrieval aims to align diverse modalities (e.g., text, images, or their combinations) within a shared embedding space for semantic similarity matching. Early works [42, 28, 56] leveraged large-scale contrastive learning to align image-text pairs, enabling zero-shot classification and cross-modal retrieval, but struggled with complex multimodal inputs. Recent advancements have addressed these limitations through innovative approaches based on MLLMs [22, 51, 24, 25]. LamRA [36] adapts MLLMs into retrievers using lightweight LoRA modules with joint pointwise and listwise reranking; MM-Embed [30] improves alignment via MLLM fine-tuning and modality-aware hard negative mining; GME [62] develops a unified multimodal retrieval framework leveraging MLLMs with synthetic data augmentation; and UniME [15] combines textual knowledge distillation from LLM teachers with hard negative-enhanced instruction tuning, achieving robust cross-modal alignment for diverse retrieval tasks. While these works provide a strong foundation for UMR, the design space for MLLM-based embedding models remains underexplored. Unlike LLMs that generate outputs via autoregressive decoding, embedding models encode features of the entire input sequence, necessitating specialized strategies for feature extraction and tailored training methodologies. This work addresses these gaps through a systematic investigation of the design decisions for MLLM-based embedding models.
3 Recipe for Building U-MARVEL
In this section, we conduct a comprehensive study to identify the key factors that drive effective embedding learning for UMR through MLLMs. Following the mainstream approaches, we first establish a simple pipeline. Unless stated otherwise, we finetune the pre-trained Qwen2-VL-7B-Instruct [47] as the embedding model with LoRA [19] by contrastive learning on M-BEIR dataset [51]. Details are given in Appendix 3. We then explore three major axes of design decisions in the following subsections:
We investigate the adaptation of decoder-only MLLMs into instruction-aware embedders.
We explore how to train embedding models within the framework of contrastive learning.
We validate that the retrieval performance can be further enhanced through distillation from the recall-then-rerank paradigm.
3.1 How to Adapt MLLMs into Embedding Models?
LLMs generate output tokens in an autoregressive manner, whereas embedding models encode holistic representations of input sequences, reflecting fundamental differences in their training objectives. Adapting LLMs into embedding models thus requires specialized training strategies. In this subsection, we examine this adaption from three aspects as follows.
3.1.1 Embedding Extraction
We first investigate the method for extracting embeddings of the token sequence of multimodal queries using MLLMs. Following prior approaches, the instruction is directly concatenated with the query. Existing methods can basically be divided into two categories based on the way embeddings are extracted. (1) Last token: previous works [36, 24, 62, 22, 15] typically employ a prompt for compression. Generally, for multimodal input, the prompt is "<image><text>…<image><text> Summarize above image and sentence in one word: <emb>", where <image> and <text> are the placeholders. In this way, the feature of last token (i.e., <emb>) is used as the output embedding of the input sequence. (2) Mean token: an alternative method, adopted by MM-Embed [30], transforms the unidirectional attention mechanism into a bidirectional one and applies mean pooling to the features from the last hidden states to derive the embedding for the entire sequence.
| ID | Causal or | Last token or | Compression Prompt | Local | Global |
| Bidirectional | Mean token | ("in one word") | Avg. | Avg. | |
| 0 | Causal | Last token | ✓ | 56.6 | 54.8 |
| 1 | Bidirectional | Last token | ✓ | 55.5 | 52.6 |
| 2 | Causal | Mean token | ✓ | 33.7 | 27.6 |
| 3 | Bidirectional | Mean token | ✓ | 51.7 | 46.1 |
| 4 | Bidirectional | Mean token | ✕ | 57.2 | 55.2 |
We conducted detailed comparative experiments, as presented in Table 1. (1) The experimental results reveal that the effectiveness of the compression prompt exhibits dual dependency. On one hand, it establishes a strong coupling with the last token mechanism, as evidenced by ID-0 and ID-2, where replacing last token with mean token results in a significant performance drop of 22.9%/27.2%. On the other hand, causal attention enhances the impact of the compression prompt (ID-0 vs. ID-1) demonstrating a 1.1%/2.2% improvement over bidirectional attention. (2) Interestingly, bidirectional attention partially decouples the dependency between the compression prompt and the last token mechanism. For instance, ID-3 shows an 18%/18.5% improvement compared to ID-2. However, when combined with mean pooling, bidirectional attention remains constrained by the limitations of the compression prompt. (3) Furthermore, when compression prompt is removed from bidirectional attention with mean token mechanism (ID-4), it achieves a modest improvement of 0.6%/0.4% over ID-0, which is commonly used in the literature.
In conclusion, generating embeddings using bidirectional attention combined with mean pooling from the sequence demonstrates superior performance compared to the commonly employed approach of combining compression prompts with the last token mechanism. This difference may stem from the last token embedding being affected by recency bias, leading to an excessive reliance on the output of the final token. Notably, this finding aligns with the conclusions of NV-Embed [26], yet diverges from those presented in GME [62], offering a distinct perspective on architectural design and performance optimization.
3.1.2 Instruction Integration
Motivated by prior studies on text embeddings [26, 4], we mask out the instruction tokens during the mean pooling process, as these tokens have already influenced the output features through self-attention mechanisms. Experimental results show that ID-0 achieves a 0.1%/0.3% improvement compared to ID-1 and 2.0%/12.1% improvement compared to ID-2 in Table 3. We hypothesize that this improvement arises primarily from that: the query inherently incorporates instruction information during forward propagation thought bidirectional self-attention mechanism. Therefore, by filtering out instruction-related information and focusing on feature comparisons between the query and candidate, the approach mitigates calculation bias.
| ID | Instructions | Masking | Local Avg. | Global Avg. |
| 0 | ✓ | ✓ | 57.3 | 55.5 |
| 1 | ✓ | ✕ | 57.2 | 55.2 |
| 2 | ✕ | - | 55.3 | 43.4 |
| ID | Methods | Local Avg. | Global Avg. |
| 0 | Instruction Tuning on M-BEIR | 56.6 | 53.9 |
| 1 | [0] + text-only retrieval | 57.3 | 55.5 |
| 2 | [1] + text-image retrieval | 57.7 | 55.8 |
3.1.3 Progressive Transition
We set a baseline by vanilla instruction tuning on multimodal retrieval data from the M-BEIR dataset, presented in Table 3 under ID-0. Following prevalent approaches [36, 22], we use text-only retrieval data for pre-training, which improves performance on various retrieval tasks. Training on the NLI dataset [13] significantly enhances our model’s performance, as indicated by ID-1 in Table 3. Despite these improvements, a notable gap persists when transitioning from single-modality, single-instruction tasks to multimodal, multi-instruction tasks.
To address this gap, we propose a progressive approach incorporating an additional pre-training step using text-image retrieval data from the CC3M dataset [43]. This step further enhances performance, as demonstrated by ID-2 in Table 3. Since the LLM was originally trained with causal attention, switching to bidirectional attention degrades text-visual encoder alignment. We speculate that text-based fine-tuning strengthens the text encoder’s semantic representation, while image-text pairs further align the text and visual encoders. During experiments, we train the model with NLI using unidirectional InfoNCE loss, followed by CC3M using bidirectional InfoNCE loss, pairing each query with its correct match (positive sample) while other items in the batch serve as negative samples.
Therefore, we formally propose the progressive transition method, adopting a stepwise training strategy to adapt decoder-only MLLMs for embedding generation tasks. Following the principle of advancing from simpler to more complex tasks, this phased approach ensures a smooth transition through three key steps: (1) Text Retrieval Adaptation: Learning retrieval tasks using text-only datasets to establish foundational capabilities. (2) Cross-modal Alignment: Advancing to cross-modal retrieval tasks through text-image paired data. (3) Instruction Tuning for Multimodal Retrieval: Mastering instruction-guided multimodal retrieval tasks with comprehensive training on multimodal datasets.
3.2 How to Train MLLM-based Embedders by InfoNCE?
In this subsection, we investigate how MLLM-based embedding models should be trained under the contrastive learning framework, following the paradigm adopted by most existing methods. Despite its simplicity, we identify that several overlooked factors exert a critical impact on the final performance, including interactions among the training parameters of InfoNCE and the control of noise in hard negative mining during the continual training process.
3.2.1 Interactions Among Large Batch Size, Learning Rate and Temperature
| ID | Batch Size | Temp () | LR () | Local Avg. |
| 0 | 480 | 0.05 (fixed) | 57.2 | |
| 1 | 1920 | 0.05 (fixed) | 57.4 | |
| 2 | 1920 | 0.05 (fixed) | 58.3 | |
| 3 | 3840 | 0.05 (fixed) | 58.5 | |
| 4 | 3840 | 0.05 (fixed) | 58.9 | |
| 5 | 5760 | 0.05 (fixed) | 58.7 | |
| 6 | 7680 | 0.05 (fixed) | 58.9 | |
| 7 | 3840 | 0.05 (learnable) | 60.1 | |
| 8 | 7680 | 0.05 (learnable) | 60.3 |
Our experimental results dispute the widely held belief that "increasing batch size directly improves performance", indicating that this assumption requires closer examination. By analyzing the results presented in Table 4, we identify three key findings as follows. (1) Experiments with ID-0, ID-2, and ID-4 indicate that increasing the batch size does improve model performance; however, the performance gains tend to plateau as the batch size continues to grow, as observed in comparisons between ID-5 and ID-6. (2) Consistent with prior research [14] that highlights the necessity of scaling the learning rate (lr) alongside batch size, our contrastive learning experiments reveal similar trends. Specifically, simply increasing the batch size without adjusting the learning rate results in marginal improvements (ID-0 vs. ID-1). Optimal learning rate scaling approximately follows the square root of the batch size increase ratio, as demonstrated in comparisons between ID-0 and ID-2. (3) Our investigation into the effects of temperature parameters reveals that employing a learnable temperature significantly enhances performance. This is evident from comparing ID-4 with ID-7 and ID-6 with ID-8. While existing methods such as LamRA [36] and MM-Embed [30] utilize a fixed temperature parameter, we argue that the learnable temperature parameter in contrastive learning dynamically adjusts the sharpness of the probability distribution, enabling the model to optimize this critical hyperparameter during training. By adapting to different training stages, dataset characteristics, and representation quality, it simplifies hyperparameter tuning, improves gradient properties, balances competing objectives, and enhances overall performance. This innovation makes contrastive learning more robust and effective across diverse applications.
It is important to note that all experiments were conducted within a single training epoch, meaning that larger batch sizes corresponded to fewer training steps. Furthermore, increasing the number of epochs provided negligible additional benefits when hyperparameters were appropriately configured.
3.2.2 Continual Training with Hard Negative Mining
The initial random negative sampling yields basic discriminative ability but struggles with challenging false positives: negatives that exhibit high semantic similarity (measured by similarity scores) yet are incorrect matches. Recent studies have employed hard negative mining [30, 25] or hard negative reweighting in loss computation [17, 41, 64] to address this issue. In this work, we explore related settings. However, as shown in Table 5 for ID-0, directly selecting the top- hard negatives proved to be excessively challenging, resulting in complete failure. Additionally, incorporating in-batch negatives results in a performance decline compared to the baseline in this table, as observed by ID-1. We hypothesize that some hard negatives are false negatives, which can adversely affect model training. To address this issue, we propose a straightforward filtering method that eliminates hard negatives with scores exceeding a predefined threshold. This approach led to a significant improvement in model performance, as demonstrated by ID-2.
We apply a hard negative mining strategy as follows: (1) Feature Extraction: Using the model from the previous training stage, we extract features for all queries and candidates, compute similarity scores (excluding positive matches), and rank candidates to identify the most challenging negatives per query; (2) Filtered Hard Negative: For each query, we filter out negative samples that exceed a predefined threshold (considered as false negatives) and then select the top- negative samples as hard negatives; (3) Balanced Training: To prevent model convergence difficulties from excessive hard negatives and to accelerate training, we select a suitable -value and mix these hard negatives with other in-batch negatives. We then perform continual finetuning by InfoNCE.
| ID | Methods | Local Avg. | Global Avg. |
| Progressive Transition (baseline) | in-batch neg | 60.6 | 58.7 |
| 0 | only top- hard neg | failed | failed |
| 1 | in-batch neg and top- hard neg | 57.4 | 55.4 |
| 2 | in-batch neg and filtered top- hard neg | 61.7 | 59.9 |
3.3 Will Reranker Distillation Improve Performance?
Some studies [36, 30] further refine retrieval results by applying a reranking step on top of the recall model. The reranker can be either a trained model [36] or a zero-shot reranker [30] based on MLLMs. While this recall-then-rerank strategy often improves retrieval performance, it introduces additional inference latency and increases system complexity. In this work, we exploit to distill this cascaded pipeline into a single unified model, thereby simplifying the system while maintaining high performance.
Inspired by [36], we train a generative reranker based on the decoder-only MLLM architecture. The training data is prepared by: (1) extracting embeddings for all queries and candidates via the recall model from Sec. 3.2.2, then (2) constructing pairs for each query consisting of a positive match and the top 50 most challenging negatives. The model is prompted to output "YES" for positives and "NO" for negatives, optimized with next-token prediction loss to maximize the likelihood of correct response.
After training the reranker, we employ it to refine the ranking results from the recall model, further improving the retrieval performance. Specifically, the recall model from Sec. 3.2.2 first retrieves the top- candidates, then each (query, candidate) pair is processed by the reranker sequentially. The reranker is prompted to judge each candidate with a "YES" or "NO" response, where the probability of "YES" serves as the reranker’s score. To further enhance overall performance, we combine the scores from both models through linear interpolation: , where becomes the final relevance score, and is a weight hyperparameter. This completes our recall-then-rerank pipeline, producing optimized query-candidate matching scores.
We distill this entire pipeline as a teacher model into a single student model, where the combined recall-then-rerank knowledge is compressed into one unified model. Specifically, the recall model is continue trained using contrastive learning with positive and top- negative samples, optimized via KL divergence:
where represents the multi-model ensemble scores and the single model’s predictions. The scores are softmax-normalized before computing the KL divergence. The inference time complexity of the model is , making full-scale distillation computationally expensive. Notably, our distillation approach achieves significant performance improvements by distilling on only of the data, demonstrating the remarkable efficiency and effectiveness of our distillation strategy.
| Methods | Local Avg. | Global Avg. |
| Hard Negative Mining (baseline) | 61.7 | 59.9 |
| Reranker | 64.5 | 61.7 |
| Distillation | 63.2 | 60.7 |
| Continue-hard | 62.2 | 60.0 |
As shown in Table 6, the "Reranker" represents the model we aim to distill, and the "Distillation" method achieves an improvement of 1.5%/0.8% over the baseline. However, "Distillation" model is built upon baseline model, with its training data generated by re-ranking the top- retrieval results from baseline model by using a "Reranker". As a result, the training data correspond to hard negative samples for the baseline model. To investigate whether the performance improvement of "Distillation" model is driven by or the effect of knowledge distillation or the introduction of hard negative samples, we designed experiment "Continue-hard" for validation. Unlike the "Distillation" stage, the "Continue-hard" model continues training by treating the training data as hard negatives. The "Continue-hard" method demonstrates an improvement of 0.5%/0.1% over the baseline; however, it exhibits a decline of 1.0%/0.7% when compared to the "Distillation" approach. It validates that the performance improvement is clearly attributable to the distillation process.
3.4 U-MARVEL Framework
Motivated by the findings presented above, we introduce a unified framework, term U-MARVEL (Universal MultimodAl RetrieVal via Embedding Learning), which consists of the following three stages: (1) Progressive transition: the model is progressively fine-tuned on the increasing complexity levels of the retrieval data, which enables the model to gradually adapt to retrieval tasks. (2) Hard Negative Mining: building on Progressive transition, the model’s retrieval performance is further enhanced through hard negative mining, accompanied by the filtering of false negatives. (3) Distillation: we perform knowledge distillation on the reranker model using the sampled dataset sequentially. Detailed experimental settings and results for U-MARVEL are provided in Table 12 in Appendix 3 and Table 13 in Appendix D, respectively.
4 Comparisons with State-of-the-art Methods
In this section, we compare U-MARVEL with state-of-the-art methods in both supervised and unsupervised settings.
4.1 M-BEIR Evaluation
We first compare U-MARVEL with competitors on M-BEIR benchmark [51] in both local and global pool settings, details of this dataset can be found in Appendix A. For fair comparison, we also present results of the recall-then-rerank cascade pipeline, denoted as , similar to that used in LamRA [36], details seen in Appendix 3. As shown in Table 7, our primary comparisons are conducted under the local pool setting, while detailed results for the global setting are provided in the Table 14 of Appendix D.
| Task | Dataset | Metric | Single Model | +Rerank | ||||||
CLIP-L [42] |
SigLIP [58] |
[51] |
[51] |
LamRA-Ret [36] |
U-MARVEL |
LamRA [36] |
| VisualNews | R@5 | 43.3 | 30.1 | 23.4 | 42.6 | 41.6 | 47.3 | 48 | 49.4 | |
| MSCOCO | R@5 | 61.1 | 75.7 | 79.7 | 81.1 | 81.5 | 84.4 | 85.2 | 85.6 | |
| Fashion200K | R@10 | 6.6 | 36.5 | 26.1 | 18 | 28.7 | 33.6 | 32.9 | 34.2 | |
| WebQA | R@5 | 36.2 | 39.8 | 80 | 84.7 | 86 | 97.1 | 96.7 | 98.5 | |
| EDIS | R@5 | 43.3 | 27 | 50.9 | 59.4 | 62.6 | 78.8 | 75.8 | 81.4 | |
| WebQA | R@5 | 45.1 | 43.5 | 79.8 | 78.7 | 81.2 | 88.5 | 87.7 | 89.4 | |
| VisualNews | R@5 | 41.3 | 30.8 | 22.8 | 43.1 | 39.6 | 47.3 | 48.6 | 50.5 | |
| MSCOCO | R@5 | 79 | 88.2 | 89.9 | 92.3 | 90.6 | 93.5 | 92.3 | 88.4 | |
| Fashion200K | R@10 | 7.7 | 34.2 | 28.9 | 18.3 | 30.4 | 35.1 | 36.1 | 37.7 | |
| NIGHTS | R@5 | 26.1 | 28.9 | 33 | 32 | 32.1 | 34.2 | 33.5 | 34.7 | |
| OVEN | R@5 | 24.2 | 29.7 | 41 | 45.5 | 54.1 | 62.5 | 59.2 | 63.7 | |
| InfoSeek | R@5 | 20.5 | 25.1 | 22.4 | 27.9 | 52.1 | 58.3 | 64.1 | 62.9 | |
| FashionIQ | R@10 | 7 | 14.4 | 29.2 | 24.4 | 33.2 | 36.4 | 37.8 | 38.2 | |
| CIRR | R@5 | 13.2 | 22.7 | 52.2 | 44.6 | 53.1 | 60.7 | 63.3 | 63.2 | |
| OVEN | R@5 | 38.8 | 41.7 | 55.8 | 67.6 | 76.2 | 79.4 | 79.2 | 80.8 | |
| InfoSeek | R@5 | 26.4 | 27.4 | 33 | 48.9 | 63.3 | 74.7 | 78.3 | 78.9 | |
| Avg. | 32.5 | 37.2 | 46.8 | 50.6 | 56.6 | 63.2 | 63.7 | 64.8 | ||
The proposed U-MARVEL framework establishes new state-of-the-art performance across both single-model architectures and recall-then-rerank approaches. Specifically: (1) U-MARVEL achieves superior results in 15 out of 16 retrieval tasks. On average, U-MARVEL outperforms models UniIR [51], MM-Embed [30] and LamRA-Ret [36], and demonstrates 6.6%/5.8%(local/global setting) improvement over LamRA-Ret, the previous state-of-the-art. (2) also outperforms LamRA [36] across the majority of tasks and achieves an average improvement of 1.1%/0.4%(local/global setting). (3) While LamRA requiring a two-stage recall-then-rerank inference process, U-MARVEL achieves comparable performance with only a single-stage inference. This breakthrough means that our training scheme has achieved the ranking quality that traditionally requires a cascade architecture on the single encoder model, achieved a balance between efficiency and effectiveness.
4.2 Zero-shot Evaluation
We evaluate the performance of U-MARVEL with several unseen datasets in a zero-shot manner. Details regarding the datasets and evaluation metrics are provided in Appendix B. As presented in Table 8, U-MARVEL achieves state-of-the-art performance on 6 out of 10 tasks, highlighting its exceptional zero-shot capabilities compared to leading methods such as VLM2Vec [24], UniME [15], and LamRA [36]. Furthermore, we evaluate U-MARVEL on the zero-shot text-to-video retrieval task. The results on the MSVD and MSR-VTT datasets, summarized in Table 9, demonstrate that U-MARVEL outperforms VLM2Vec [24], LamRA [36], and LLaVE-7B [25]. However, it is worth noting that exhibits degraded performance, which may be attributed to fail to model temporal contexts, such as action sequence progression.
| Task | Dataset | Metric | Single Model | +Rerank | ||||||||
CLIP-L [42] |
[51] |
E5-V [22] |
MagicLens-L [60] |
VLM2Vec [24] |
UniME [15] |
LamRA-Ret [36] |
U-MARVEL |
LamRA [36] |
| ShareGPT4V | R@1 | 84 | 85.8 | 86.7 | 85.5 | 90.7 | 97.2 | 93.3 | 96.4 | 97.9 | 97.8 | |
| R@1 | 52.8 | 75 | 84 | 59.3 | 90.8 | 95.9 | 95.1 | 96.7 | 98.8 | 97.7 | ||
| Flickr | R@1 | 67.3 | 78.7 | 79.5 | 72.5 | 76.0 | 81.9 | 82.8 | 85.4 | 88.1 | 88.5 | |
| ShareGPT4V | R@1 | 81.8 | 84.1 | 84 | 60.9 | 85.8 | 93.9 | 88.1 | 97.2 | 96.5 | 98.9 | |
| R@1 | 68.7 | 78.4 | 82.4 | 24.2 | 84.7 | 95.2 | 94.3 | 93.5 | 98 | 98.2 | ||
| Flickr | R@1 | 87.2 | 94.2 | 88.2 | 84.6 | 90.6 | 93.4 | 92.7 | 93.3 | 97.6 | 95.1 | |
| MAP@5 | 4 | 12.5 | 24.8 | 29.6 | - | - | 33.2 | 36.2 | 42.8 | 46.0 | ||
| R@1 | 13.3 | 16.8 | 18.5 | 16.3 | - | - | 18.9 | 19.1 | 24.8 | 22.6 | ||
| R@1 | 23.7 | 26.8 | 54.6 | 28 | - | - | 62.8 | 70.3 | 70.9 | 77.6 | ||
| R@5 | 17.7 | 39.4 | 19.2 | 22.6 | - | - | 60.9 | 65.7 | 63.9 | 66.3 | ||
| ITM | CC-Neg | Acc. | 66.7 | 79.9 | 83.2 | 62.7 | - | - | 79.6 | 84.5 | 85.9 | 86.1 |
| Acc. | 73 | 80.3 | 84.7 | 75.9 | 79.5 | 85.0 | 85.8 | 87.9 | 93.5 | 93.4 | ||
| * indicates that the images in these datasets are sourced from COCO or FashionIQ. | ||||||||||||
| Model | MSR-VTT | MSVD | ||||
| R@1 | R@5 | R@10 | R@1 | R@5 | R@10 | |
| Zero-shot (finetuned with text-video data) | ||||||
| InternVideo [50] | 40.0 | 65.3 | 74.1 | 43.4 | 69.9 | 79.1 |
| ViCLIP [48] | 42.4 | - | - | 49.1 | - | - |
| UMT-L [29] | 42.6 | 64.4 | 73.1 | 49.9 | 77.7 | 85.3 |
| InternVideo2s2-6B [49] | 55.9 | 78.3 | 85.1 | 59.3 | 84.4 | 89.6 |
| Zero-shot (finetuned only with text-image data) | ||||||
| VLM2Vec [24] | 43.5 | 69.3 | 78.9 | 49.5 | 77.5 | 85.7 |
| LamRA [36] | 44.7 | 68.6 | 78.6 | 52.4 | 79.8 | 87.0 |
| LLaVE-7B [25] | 46.8 | 71.1 | 80.0 | 52.9 | 80.1 | 87.0 |
| U-MARVEL | 47.2 | 72.0 | 80.2 | 54.6 | 80.9 | 87.7 |
| 47.5 | 69.7 | 78.5 | 53.1 | 79.9 | 87.1 | |
5 Conclusion and Limitations
This paper systematically explores the design principles for building high-performing universal multimodal retrievers with MLLMs. Through a comprehensive study, we identified key factors that significantly impact retrieval performance, including embedding generation strategies, progressive adaptation techniques, and training recipes such as hard negative mining and re-ranker distillation. Based on these insights, we introduce U-MARVEL, a unified framework that achieves state-of-the-art performance on the M-BEIR benchmark in supervised settings and demonstrates strong zero-shot generalization across diverse tasks. The results highlight the generalization potential of U-MARVEL and its ability to address the complex requirements of real-world retrieval scenarios. We believe that our study contributes valuable knowledge to the research community and paves the way for future advancements in universal multimodal retrieval.
Despite the promising results achieved by U-MARVEL, there remain several limitations that can be further improved. First, it primarily focuses on text and image modalities, leaving the inclusion of other modalities (e.g., audio) for future work. Second, while effective in zero-shot scenarios, its integration with LLMs in retrieval-augmented generation (RAG) applications remains underexplored. Lastly, experiments were limited to 7B-sized models due to submission time constraints; future work will extend support models of various sizes, enabling flexible trade-offs between performance and cost.
Appendix
Appendix A M-BEIR Dataset & Metrics
Datasets:
We evaluate the retrieval performance of our method with the M-BEIR [51] benchmark. This dataset consist of 8 possible multimodal retrieval tasks involving texts and images. It is constructed based on 10 different datasets from 4 domains. Moreover, according to the datasets and task types (i.e., the modalities of queries and candidates), the multi-modal retrieval tasks are further subdivided into 16 types.
Regarding the structure of the dataset, the M-BEIR consists of three key components: instructions, query sets, and candidate pools. Each task has its independent query set and candidate pool, and the query set and candidate pool of each task contain only a single modality. Additionally, each task is equipped with 4 instructions, which clearly specify the target modality of the task. Regarding the scale of the dataset, the training set contains 1.33 million queries and 1.93 million candidates, while the test set contains 0.19 million queries and 5.6 million candidates. In addition, besides the local query sets and candidate pools for each task, both the training set and the test set also have global query sets and global candidate pool, that is, the query sets and candidate pools of the 16 tasks are combined. Table 10 provides a detailed presentation of the information about the M-BEIR dataset.
| Task | Dataset | Instruction (shown 1 out of 4) | Domain | Train | Dev | Test | Pool |
| VisualNews [33] | Identify news-related image match with the description | News | 99K | 20K | 20K | 542K | |
| MSCOCO [31] | Find an everyday image match with caption | Misc. | 100K | 24.8K | 24.8K | 5K | |
| Fashion200K [16] | Based on fashion description, retrieve matched image | Fashion | 15K | 1.7K | 1.7K | 201K | |
| WebQA [5] | Find an paragraph from Wikipedia to answer the question | Wiki | 16K | 1.7K | 2.4K | 544K | |
| EDIS [35] | Find a news image matching with the caption | News | 26K | 3.2K | 3.2K | 1M | |
| WebQA [5] | Find a Wiki image that answer the question | Wiki | 17K | 1.7K | 2.5K | 403K | |
| VisualNews [33] | Provide a news-related caption for the displayed image | News | 100K | 20K | 20K | 537K | |
| MSCOCO [31] | Find a caption describe the an image | Misc. | 113K | 5K | 5K | 25K | |
| Fashion200K [16] | Find a description for the fashion item in the image | Fashion | 15K | 4.8K | 4.8K | 61K | |
| NIGHTS [12] | Find an image that is identical to the given image | Misc. | 16K | 2K | 2K | 40K | |
| OVEN [20] | Retrieve a Wiki text that answer the given query about the image | Wiki | 150K | 50K | 50K | 676K | |
| InfoSeek [7] | Find an article that answers the given question about the image | Wiki | 141K | 11K | 11K | 611K | |
| FashionIQ [52] | Find an image to match the fashion image and style note | Fashion | 16K | 2K | 6K | 74K | |
| CIRR [38] | I’m looking for a similar everyday image with the described changes | Misc. | 26K | 2K | 4K | 21K | |
| OVEN [20] | Find a Wiki image-text pair to answer a question regarding an image | Wiki | 157K | 14.7K | 14.7K | 335K | |
| InfoSeek [7] | Find a Wiki image-text pair to answers my question about this image | Wiki | 143K | 17.6K | 17.6K | 481K | |
| 10 datasets | 64 instructions | 4 domains | 1.1M | 182K | 190K | 5.6M |
Metrics:
Following the evaluation protocols by M-BEIR benchmark, we use the Recall@10 (R@10) for Fashion200K [16] and FashionIQ [52], and the Recall@5 (R@5) for the remaining datasets. We report the average score of all test queries as the criterion for measuring the retrieval accuracy.
The evaluation has two kinds of settings: Local and Global pool settings. The local setting indicates that each task conducts retrieval operations within its own candidate pool and the retrieved results will not have modality errors. The global setting means that the retrieval is carried out within the global candidate pool formed by combining all task candidate pools. In this case, modality errors will occur in the retrieved results.
Appendix B Zero-shot Datasets & Metrics
Datasets:
We have examined the zero-shot capabilities of U-MARVEL in retrieval tasks, specifically focusing on image-text cross-modal retrieval and text-video retrieval. For tasks of cross-modal retrieval between images and texts, we use the following datasets: ShareGPT4V [21], Urban-1K [59], CIRCO [3], Flickr [40], GeneCIS [46], Visual Dialog [11], Multi-round FashionIQ [57], CC-Neg [44], and Sugar-Crepe [18]. For text-video retrieval, the datasets used are MSR-VTT [53] and MSVD [6]. We present the details of the Unseen Dataset in Table 11. Although some of them are actually adapted from MSCOCO or FashionIQ, we still treat these datasets as unseen datasets due to the significant differences in their captions or query formats, following [36]. For instance, the captions in Urban1K consist of extended captions generated by GPT-4V [21], while the query format of CIRCO combines a reference image with a relative caption. These differences create a substantial disparity compared to the original COCO [31] dataset. Table 11 provides a detailed presentation of the information about the zero-shot datasets.
Metrics.
The first row of Table 8 indicates the retrieval task type: for text queries, for image queries, for dialog queries, for multi-interleaved image-text queries, for text candidates, for image candidates, and ITM for the Image-Text Matching task.
| Benchmark | Zero-shot | # Queries | # Candidates |
| ShareGPT4V [21] | 1K | 1K | |
| Urban-1K [59] | 1K | 1K | |
| Flickr30K [40] | 1K | 5K | |
| CIRCO [3] | 800 | 120K | |
| GeneCIS [46] | 8K | 10 - 15 | |
| Visual Dialog [11] | 2K | 2K | |
| Multi-round FashionIQ [57] | 2.4K | 6.2K | |
| CC-Neg [44] | 40K | 2 | |
| Sugar-Crepe [18] | 7.5K | 2 | |
| MSR-VTT [53] | 670 | 27K | |
| MSVD [6] | 1K | 1K |
Appendix C Implementation Details
We select Qwen2-VL-7B-Instruct [47] as the pretrained MLLM backbone to instantiate both of the retriever and the reranker. Throughout all experiments, we froze the visual side and employed LoRA [19] to fine-tune the LLM component of the model using Ascend-910B GPUs. For embedding models, the InfoNCE loss [39] is selected as our training objective:
where , , represent the embeddings of the query, the positive candidate, and any sample from the candicate set, respectively. is the temperature parameter. means the cosine similarity.
For reranking models, we train point-wise rerankers to classify whether a pair of query and candidate items is relevant. The loss function for the reranking model is formally defined as follows:
where if the candidate belongs to the set of positive candidates, and otherwise. Detailed training configuration are shown in Table 12.
| Training Stage | Temp () | LR () | Batch Size | GPU number | LoRA para. | epoch number |
| Progressive Transition (stage-1) | 0.05 (fixed) | 576 | 8 | rank(64)(128) | 2 | |
| Progressive Transition (stage-2) | 0.05 (fixed) | 720 | 16 | rank(128)(256) | 1 | |
| Progressive Transition (stage-3) | 0.050.0206* | 3840 | 64 | rank(128)(256) | 1 | |
| Hard-neg mining | 0.02060.0188* | 1408 | 64 | rank(128)(256) | 1 | |
| Distillation | 0.01880.0204* | 50 | 64 | rank(128)(256) | 1 | |
| Reranker Training | - | 128 | 64 | rank(128)(256) | 1 | |
| *Learnable temperature parameter | ||||||
We use the U-MARVEL model as the initial retrieval system and employ its Top-100 candidate results to train a point-wise reranking model following the methodology presented in Figure 1(b) and Eq. 3 as the loss function. Subsequently, the embedding similarity scores produced by the U-MARVEL model are combined with the reranking scores from the point-wise reranking model through a weighted fusion mechanism to compute the final ranking score:
where is a tunable weighting hyperparameter. Based on this final score, the candidate set is reranked to generate the final ranked list. We refer to this unified framework as . The detailed training configurations are provided in Table 12.
Appendix D Supplemental Experimental Results
Effectiveness of three key stages in U-MARVEL.
We present the retrieval performance of each stage in U-MARVEL in Table 13, which provides a detailed ablation study of our framework. The results clearly demonstrate the effectiveness of the proposed design, showcasing significant performance improvements achieved through the progressive transition mechanism, hard negative mining, and reranker distillation. These components collectively contribute to the robustness and efficiency of U-MARVEL, highlighting the importance of each stage in enhancing retrieval capabilities and validating the overall framework design.
Comparison of M-BEIR (global).
Table 14 presents a comprehensive comparison of our proposed U-MARVEL framework with state-of-the-art methods on the M-BEIR benchmark under the global pooling setting. The results demonstrate that, within the single-model category, U-MARVEL significantly outperforms leading approaches such as MM-Embed [30] and LamRA-Ret [36], by a large margin. Furthermore, when integrated into a recall-then-rerank pipeline, the enhanced variant achieves the best overall performance, further solidifying the effectiveness of our framework. These findings highlight the versatility and superiority of U-MARVEL across different retrieval paradigms, showcasing its potential for advancing state-of-the-art retrieval systems.
Held-out experiments on M-BEIR.
To evaluate the cross-task generalization ability of the proposed method, we trained the model on the five tasks, reserving the remaining three tasks as held-out tasks for evaluation. As presented in Table D, our method achieves an average improvement of 2.4%/1.6% across the three held-out tasks compared with SoTA approaches.
| Task | Dataset | Local | Global | ||||
| Progressive | Hard-neg | Distillation | Progressive | Hard-neg | Distillation | ||
| transition | mining | transition | mining | ||||
| VisualNews | 46.4 | 47.5 | 47.3 | 46.3 | 47.5 | 47.2 | |
| MSCOCO | 84.0 | 83.6 | 84.8 | 77.3 | 79.4 | 72.8 | |
| Fashion200K | 34.4 | 35.6 | 33.6 | 34.3 | 35.4 | 33.3 | |
| WebQA | 92.4 | 95.2 | 97.1 | 92.2 | 95.0 | 96.7 | |
| EDIS | 65.6 | 72.9 | 78.8 | 65.4 | 72.9 | 78.7 | |
| WebQA | 84.6 | 86.9 | 88.5 | 83.8 | 86.6 | 87.7 | |
| VisualNews | 45.7 | 47.9 | 47.3 | 45.4 | 47.8 | 47.2 | |
| MSCOCO | 93.3 | 93.1 | 93.5 | 93.2 | 93.1 | 93.5 | |
| Fashion200K | 35.4 | 35.6 | 35.1 | 35.3 | 35.6 | 34.9 | |
| NIGHTS | 31.4 | 33.0 | 34.2 | 31.4 | 33.0 | 34.0 | |
| OVEN | 56.5 | 58.9 | 62.5 | 51.3 | 52.2 | 58.3 | |
| InfoSeek | 55.1 | 52.2 | 58.3 | 50.8 | 47.5 | 52.2 | |
| FashionIQ | 35.4 | 34.8 | 36.4 | 35.3 | 34.7 | 36.0 | |
| CIRR | 58.7 | 57.2 | 60.7 | 55.9 | 55.2 | 56.0 | |
| OVEN | 78.0 | 80.8 | 79.4 | 71.9 | 73.8 | 73.1 | |
| InfoSeek | 72.7 | 72.3 | 74.7 | 68.5 | 68.7 | 69.2 | |
| Avg. | 60.6 | 61.7 | 63.2 | 58.7 | 59.9 | 60.7 | |
| Task | Dataset | Metric | Single Model | +Rerank | |||||
[51] |
[51] |
MM-Embed [30] |
LamRA-Ret [36] |
U-MARVEL |
LamRA [36] |
| VisualNews | R@5 | 23 | 42.6 | 41 | 41.3 | 47.2 | 46.9 | 48.8 | |
| MSCOCO | R@5 | 75.6 | 77.9 | 71.3 | 75.4 | 72.8 | 78 | 70.1 | |
| Fashion200K | R@10 | 25.4 | 17.8 | 17.1 | 28.7 | 33.3 | 32.5 | 33.8 | |
| WebQA | R@5 | 79.5 | 84.7 | 95.9 | 85.8 | 96.7 | 96.5 | 98.3 | |
| EDIS | R@5 | 50.3 | 59.4 | 68.8 | 62.5 | 78.7 | 74.4 | 80.8 | |
| WebQA | R@5 | 79.7 | 78.8 | 85 | 81 | 87.7 | 87.1 | 88.3 | |
| VisualNews | R@5 | 21.1 | 42.8 | 41.3 | 39.3 | 47.2 | 47.6 | 49.8 | |
| MSCOCO | R@5 | 88.8 | 92.3 | 90.1 | 90.4 | 93.5 | 92.4 | 86.0 | |
| Fashion200K | R@10 | 27.6 | 17.9 | 18.4 | 30.4 | 34.9 | 36.6 | 36.8 | |
| NIGHTS | R@5 | 33 | 32 | 32.4 | 32.1 | 34.0 | 34.2 | 34.8 | |
| OVEN | R@5 | 38.7 | 39.2 | 42.1 | 48.4 | 58.3 | 54 | 58.7 | |
| InfoSeek | R@5 | 19.7 | 24 | 42.3 | 48.7 | 52.2 | 58.7 | 56.9 | |
| FashionIQ | R@10 | 28.5 | 24.3 | 25.7 | 33.1 | 36.0 | 37.4 | 37.4 | |
| CIRR | R@5 | 51.4 | 43.9 | 50 | 50.5 | 56.0 | 59.7 | 58.4 | |
| OVEN | R@5 | 57.8 | 60.2 | 64.1 | 70 | 73.1 | 72.6 | 74.9 | |
| InfoSeek | R@5 | 27.7 | 44.6 | 57.7 | 60 | 69.2 | 74 | 73.8 | |
| Avg. | 45.5 | 48.9 | 52.7 | 54.9 | 60.7 | 61.4 | 61.8 | ||
| Methods | NIGHTS | OVEN | InfoSeek | OVEN | InfoSeek | Local Avg. |
| R@5 | R@5 | R@5 | R@5 | R@5 | ||
| Supervised | ||||||
| UniIR-BLIP [51] | 33.0 | 41.0 | 22.4 | 55.8 | 33.0 | 37.0 |
| UniIR-CLIP [51] | 32.0 | 45.5 | 27.9 | 67.6 | 48.9 | 44.4 |
| Zero-shot | ||||||
| LamRA-Ret* [36] | 27.2 | 44.7 | 44.0 | 62.8 | 49.5 | 45.6 |
| U-MARVEL* | 31.5 | 44.3 | 49.2 | 62.6 | 52.6 | 48.0 |
| LamRA* [36] | 29.2 | 46.9 | 54.2 | 65.1 | 59.1 | 50.9 |
| * | 32.1 | 49.0 | 55.3 | 64.9 | 61.2 | 52.5 |